# Testing Hypotheses of Covariate Effects on Topics of Discourse

### v1

@@ 6 Conclusion
    In natural sciences, measurements are selected based on available equipment and often hypotheses must be assessed using observable proxies of interest. Statistician are given data and analysis performed in accordance with the experimental design. In data science, the division between the data collector and data analyst is blurred. The data scientist is tasked with incorporating domain expertise into the pre-processing of the dataset. Data curation and cleaning decisions require care and will impact results. When exploring regional differences in beers, all mentions of geography had to be removed from the text to prevent mention of Ontario in the text as being the driver for the Ontario covariate effect. With text especially, data decisions impact the outcome similar to how tuning experimental equipment in lab sciences must be conditioned upon when analyzing data.
@@
    The selection of topics as dependent variables requires care in how hypotheses are crafted to avoid the perils of using the same dataset for hypothesis generation and testing (Egami et al., 2022). The Beer hypotheses were constructed from external insights and interests around flavours, however determining anchor based topics that measure attributes of interest requires examination of $\Phi$, which in turn is based on the data. Reproducing the tests on a new independent dataset could be performed by bringing $\Phi$ or by carrying forward anchors. Even then, terminological drift suggests that application of the topic metric requires consideration of the utility of $\Phi$ within the new scientific context.
@@
    Increased discourse about a flavour topic may or may not represent actual difference in flavour probability. Unstructured text should not be used as a replacement for scientific measurement, but should be considered as an association to be used along side domain expertise to refine hypotheses for more rigorous testing. As much as it is important to use data filtered to the context of interest to avoid introducing noisy topics, additional strength of results will be provided by refining the population of writers who provide the text data. More recent advances in Large Language Models have shown promise in categorizing text, however PTMs provide a magnitude of an observable feature rather than a binary categorical presence or absence. However there is considerable potential for exploiting LLMs as part of the data pre-processing and cleaning pipelines.
    The main advantage to the BRETT approach is its stability. The NMF decomposition with anchor words is not subject to random seeds. Anchor words can be found automatically or (partially) provided by the analyst to fine tune hypotheses. The stability in BRETT allows covariate effects to be directly modelled and uses fewer approximations than STMs.

### v2

@@ 6 Discussion
    In natural sciences, measurements are selected based on available equipment and often hypotheses must be assessed using observable proxies of interest. Statistician are given data and analysis performed in accordance with the experimental design. In data science, the division between the data collector and data analyst is blurred. The data scientist is tasked with incorporating domain expertise into the pre-processing of the dataset. Data curation and cleaning decisions require care and will impact results. When exploring regional differences in beers, all mentions of geography had to be removed from the text to prevent mention of Ontario or geographic proxies as being the driver for the Ontario covariate effect. Data cleaning decisions impact the outcome similar to how tuning experimental equipment in lab sciences must be conditioned upon when analyzing data.
@@
    The selection of topics requires care in how hypotheses are crafted to avoid the perils of using the same dataset for hypothesis generation and testing (Egami et al., 2022). The Beer hypotheses were constructed from external insights and interests around flavours, however determining anchor based topics that measure attributes of interest requires examination of $\Phi$, which in turn is based on the data. The simulation study of section 4.2 suggests that using $\Phi$ from the NMF decomposition of an external independent corpus will increase the MSE relative to simply carrying forward the anchor terms and applying the decomposition on the new dataset. We expect that this will depend on the distance between word probabilities across the datasets though a thorough exploration thereof and it’s implications for experimental design is beyond the scope of this work.
    While this paper considers OLS and Beta regression models, more complex regression models could be used. Generalized Additive Models or Deep Learning for Beta distributed data are natural extensions that follow from the BRETT framework.
@@
    Increased discourse about a flavour topic may or may not represent actual difference in flavour probability. Unstructured text should not be used as a replacement for scientific measurement, but should be considered as an association to be used along side domain expertise to refine hypotheses for more rigorous testing. As much as it is important to use data filtered to the context of interest to avoid introducing noisy topics, additional strength of results will be provided by refining the population of writers who provide the text data. More recent advances in Large Language Models have shown promise in categorizing text, however PTMs provide a magnitude of an observable feature rather than a binary categorical presence or absence. There remains considerable potential for exploiting LLMs as part of the data pre-processing and cleaning pipelines. Embedding models such as BERT define lower dimensional latent word spaces that map words closer or further apart depending on their semantic meaning. As the vocabulary becomes more general, these tools are expected to provide considerable advantages in text modelling.
    In many text applications time is an important factor, and to properly consider serial correlation in discourse, this should be considered within the NMF decomposition. Non-Negative Tensor Factorization may provide a useful starting point for future work on time-topic decomposition of text.